Algorithm MyAlgorithm (A,B) Input:

Arrays A and B each storing n >= 1 integers.

Output: What is the output? (Refer to part b below)

Start: count = 0

C = 10 for i = 0 to C do

{

sum = 0 for j = 0 to n-1 do {

sum = sum + A[0]

for k = 1 to j do

sum = sum + A[k]

}

if B[i] == sum then count = count + 1

}

return count

Document a hand-run on MyAlgorithm for input arrays A = [9 2 5 1] and B = [40 29 2 57] and show the final output.
Answer #1

Initially the values of n is 4

Arrays A = [9 2 5 1] and B = [40 29 2 57]

count = 0 and C = 10

C = 10

for i = 0

sum = 0

for j = 0

sum = sum + 9 =0+9=9

for k = 1 to 0 do // condition fail

so sum is 9

for j = 1

sum = sum + 9 =9+9=18
for k = 1 to 1 do

k=1:

sum = sum + A[1] = 18 + 2 = 20

so sum is 20

for j = 2

sum = sum + 9 =20+9=29
for k = 1 to 2 do

k=1:

sum = sum + A[1] = 29 + 2 = 31

k=2:

sum = sum + A[2] = 31 + 5 = 36

so sum is 36

for j = 3

sum = sum + 9 =36+9=45
for k = 1 to 3 do

k=1:

sum = sum + A[1] = 45 + 2 = 47

k=2:

sum = sum + A[2] = 47 + 5 = 52

k=3:

sum = sum + A[3] = 52 + 1 = 53

so sum is 53

for j = 4 // condition fail

if 40 == 53 // condition fail

now i=1

Repeat the same process

Final comparisons as follows

if 29 == 53 // condition fail

if 2 == 53 // condition fail

if 57 == 53 // condition fail

Therefore the count is 0.

return 0